# From Architectures to Applications: A Review of Neural Quantum States

### v2

@@ 2.4 Graph Neural Networks
2.5 Latent Space Representations: Autoencoders
An autoencoder consists of an encoder, the latent space or bottleneck layer, and a decoder, see Fig. 6. The encoder, typically consisting of several fully connected or convolutional layers, compresses the input into just a few nodes in the bottleneck layer. The decoder subsequently generates an output based on the information in the bottleneck layer. The network parameters are optimized such that the generated output is as close to the input as possible. In a variational autoencoder [111], the encoder generates the values for mean values and variances in the bottleneck layer, and the values used as input for the decoder are then sampled from a multivariate Gaussian.
@@ 2.6.2 Transformers
A transformer model relies entirely on an attention mechanism which draws global dependencies between input and output [132]. This self-attention layer in the transformer setup generates all-to-all interactions between the sites in the system. These trainable connections can potentially represent strong connections or correlations, regardless of their position [133, 134]. The transformer first embeds the different given input elements into a unified feature space. This embedding corresponds to a linear projection, with trainable parameters, of the input elements with a dimension $d_{i}$ to elements with an embedded dimension $d_{h}$. The position of the inputs in the sequences are not explicitly modeled in the transformer, but efficiently transformed into abstract representations using positional encoding vectors that are added to the embedded input vectors [134]. Each embedded input element $\sigma_{i}$ is projected on a query vector ($q_{i}$), key vector ($k_{i}$) and a value vector ($v_{i}$) of the same dimension $d_{h}$ as the embedded input, given by:
$q_{i}=\sum_{l=1}^{d_{h}}W_{i,l}^{q}\sigma_{i,l}$ $k_{i}=\sum_{l=1}^{d_{h}}W_{i,l}^{k}\sigma_{i,l}$ $v_{i}=\sum_{l=1}^{d_{h}}W_{i,l}^{v}\sigma_{i,l},$ (15)
with the matrices $W^{q},W^{k}$ and $W^{v}$ to be the trainable weight matrices of dimension $d_{h}\crossproduct d_{h}$. The query, key and value matrices are then given by $Q=(q_{1},...,q_{N})$, $K=(k_{1},...,k_{N})$ and $V=(v_{1},...,v_{N})$. The model can be made autoregressive by using a masked self-attention layer, which allows connections to all previous elements in the sequence [132]. This autoregressive structure enables efficient exact sampling from the model [135, 128]. Then, after the masking term $M$ is added to the signal, a softmax activation function is applied to the masked dot product of the vectors $Q$ and $K$. In multi-headed attention, each query, key and value vector is mapped to $h$ vectors with trainable weight matrices, with $h$ the number of attention heads. This is indicated in Fig. 8 (middle). The complete attention formalism can be summarized by
@@
as shown in Fig. 8 (right).
Transformer quantum states can learn ground state properties of various physical systems, such as the 1D transverse field Ising model and the 1D Heisenberg $J_{1}-J_{2}$ and XYZ model [135, 134]. Comparable results to DMRG calculations have been obtained for a 1D frustrated spin model, with a relatively low number of parameters [134]. In Ref. [128], transformers have shown to be able to simulate the real-time dynamics and steady state in 1D and 2D transverse field Ising and Heisenberg models [128]. In Refs. [136, 137, 138, 139] transformers are used in the context of quantum chemistry calculations.
Small modifications of the transformer model, leading to the so-called vision transformers (ViT), inspired the use of patched transformers. This model splits the system into patches, and can be used to calculate the ground state properties of frustrated spin models [134]. A large patch size enhances the efficiency of the transformer, but on the other hand the network output dimension increases exponentially with the patch size, as it encodes the probability distribution over all possible patch states. To overcome this, large patched transformers are introduced in Ref. [133]. In this ansatz, the output of the patched transformer is passed to a patched RNN as the initial hidden state, which breaks the large inputs into smaller sub-patches, reducing the output dimension. This model has been shown to accurately capture ground state properties and phase transitions of large Rydberg systems, which can compete with quantum Monte Carlo results [133].
In Ref. [135], transformer quantum states have been used to learn ground state properties of a single system, as well as to generalize to different, unseen systems. For the latter, not only the physical degrees of freedom, but also the parameters of the Hamiltonian of the system are used as input. These parameters can be formulated as new elements that have to be passed to the embedding layer. After training the transformer quantum state for a Hamiltonian with different parameters, the transformer is able to generate the ground state for unseen Hamiltonian parameters without any additional training. Although these are with slightly larger error, more accurate results can be achieved with less training then without any a priori training.
@@ 2.7.1 Fermionic Architectures
For quantum many-body systems, mostly second quantization is used, despite some exceptions e.g. for repulsively interacting, spin-polarized fermions [147], where the authors chose to model $\psi_{0}(\bm{\sigma})\mathcal{J}(\bm{\sigma})$ in Eq. (17) by a single neural network. In the second quantization approaches used in the literature, in many cases machine learning approaches are used to enhance the expressivity of the Slater-Jastrow ansatz. This can be done by employing NNs to parameterize the Jastrow factor $\mathcal{J}$, or by using machine learning methods to dress the Slater determinant. For the latter, usually two methods are used [148]: $(i)$ The hidden fermion determinant state, where neural networks are used to replace the standard Slater determinant with a larger determinant which includes single particle orbitals from additional projected hidden fermions [149]. $(ii)$ Neural backflow transformations, which add correlation by making the single particle orbitals of the Slater determinant configuration dependent, with the respective transformation learned by a neural network [150, 151, 143, 142, 109].
One of the first examples for fermionic NQS is the RBM+PP architecture in Ref. [152], with a slightly different ansatz than Eq. (17), i.e.
@@ 2.7.2 Bosonic Architectures with Jordan Wigner Strings
The other way to simulate fermionic systems using NQS are Jordan-Wigner (JW) transformations, which is used to map the bosonic NQS to a fermionic wave function. Hence, per se bosonic architectures can be used and no special fermionic architecture is needed. The JW transformation is given by
@@
where indices refer to a one-dimensional labeling of the fermions, the $\hat{\sigma}^{\pm}$ are the spin raising and lowering operators and $\hat{c}$ fulfill the fermionic commutation relations. More precisely, for two fermions at site $i$ and $j$, exchanging these fermions yields a minus sign arising from the argument of the exponential in Eq. (19). This rule does not have to be implemented in the NQS architecture itself, but only on the level of the calculation of expectation values. For an operator $\hat{O}$ with $\langle\hat{O}\rangle=\langle O^{\mathrm{loc}}_{\bm{\theta}}(\bm{\sigma})\rangle_{\bm{\sigma}}$ (see Eq. (3)) with
@@ 2.8 Other Design Choices
A second design choice is how to incorporate symmetries in the NQS training in order to restrict the optimization space to states in the target symmetry sector, hence improving the performance [83, 123, 117, 45]. Firstly, global $U(1)$ symmetries can be imposed, see e.g. Refs. [123, 117, 158], by restricting to configurations $\bm{\sigma}$ that obey the respective symmetry, e.g. magnetization or particle number conservation. For autoregressive architectures, this is done by restricting the conditional probabilities to the targeted symmetry, see e.g. Sec. 2.6.1. For non-autoregressive architectures, the Mone Carlo updates can be chosen such that all generated configurations stay in the same symmetry sector. Second, spatial symmetries can be imposed. There are different symmetrizations used in the literature [^1] We follow Ref. [45] here. :
@@
    $\psi^{S}_{\bm{\theta}}(\bm{\sigma})=\sqrt{\mathrm{exp}\left(\frac{1}{|\mathcal{T}|}\sum_{\bm{\sigma}^{S}\in\mathcal{T}(\bm{\sigma})}2\mathrm{Re}\{\mathrm{log}\left[\psi_{\bm{\theta}}(\bm{\sigma}^{S})\right]\}\right)\cdot\mathrm{exp}\left(i\,\mathrm{arg}\left(\sum_{\bm{\sigma}^{S}\in\mathcal{T}(\bm{\sigma})}\mathrm{exp}\left(i\,\mathrm{Im}\{\mathrm{log}\left[\psi_{\bm{\theta}}(\bm{\sigma}^{S})\right]\}\right)\right)\right)}$ (23)
for symmetry operations $\mathcal{T}$ on samples $\bm{\sigma}$. Note that only the last keeps the autoregressive property intact since $\sum_{\bm{\sigma}^{S}\in\mathcal{T}(\bm{\sigma})}|\psi_{\bm{\theta}}^{S}(\bm{\sigma}^{S})|^{2}=\sum_{\bm{\sigma}^{S}\in\mathcal{T}(\bm{\sigma})}|\psi_{\bm{\theta}}(\bm{\sigma}^{S})|^{2}$. Furthermore, architectures that preserve certain symmetries explicitly can be used, such as group CNNs [99] and gauge equivariant neural Networks [159].
@@ 3.1 Ground States
For autoregressive architectures, $\langle\psi_{\bm{\theta}}|\psi_{\bm{\theta}}\rangle=1$ and hence the second term vanishes when calculating the derivative of Eq. (3). However, often $H_{\mathrm{loc}}(\bm{\sigma})$ is replaced by the covariate $\left(H_{\mathrm{loc}}(\bm{\sigma})-\langle H_{\mathrm{loc}}\rangle\right)$ to reduce the variance of the gradients [173, 117], leading to the same expression as Eq. (25). Another approach is to (pre-)train the NQS with experimental or numerical data, see Sec. 3.6.1.
@@
Further improvement can be achieved using complementary optimization methods: Firstly, the NQS can be pretrained with external data, see Sec. 3.6. Second, the energy resulting from the VMC optimization can be improved by applying a few Lanczos steps after optimizing the network parameters using the techniques described above [193]. Applying the Hamiltonian in the Lanczos algorithm to obtain the next Krylov vector corresponds to minimizing the infidelity to the corresponding state and thus necessitates a separate training, rendering typically only very few Krylov vectors accessible. Similar in spirit to Lanczos algorithms, power methods can be used to find the ground states of (gapped) Hamiltonians [194].
@@ 3.2 Excited States
The former usually rely on the same VMC scheme, but with a restriction to the targeted symmetry sector implemented in the wave function or the training loss [54, 35, 79, 157, 199]. This enables e.g. the calculation of quasiparticle dispersions from NQS [54, 35, 79, 74]. For example, in Ref. [35], specific momentum $\bm{k}_{\mathrm{target}}$ sectors are targeted by adding a mean square error between $\bm{k}_{\mathrm{target}}$ and $\bm{k}_{\mathrm{NQS}}$, see Fig. 10a. Hereby, the momentum of the NQS is given by

### v3

@@ 2.5 Autoregressive Networks
Some architectures like CNNs and transformers (see Sec. 2.6) can be made autoregressive by masking out future inputs $\sigma_{i},\dots\sigma_{N}$ for the $i$-th input vector [34, 110, 98, 101, 103, 110]. In the following, we discuss a network architecture to which the autoregressive property is inherent due to their recurrent structure: recurrent neural networks.
@@ 2.6 Transformers
A transformer model relies entirely on an attention mechanism which draws global dependencies between input and output [148]. This self-attention layer in the transformer setup generates all-to-all interactions between the sites in the system. These trainable connections can potentially represent strong connections or correlations, regardless of their position [101, 97]. The transformer first embeds the different given input elements into a unified feature space. This embedding corresponds to a linear projection, with trainable parameters, of the input elements with a dimension $d_{i}$ to elements with an embedded dimension $d_{h}$. The position of the inputs in the sequences are not explicitly modeled in the transformer, but efficiently transformed into abstract representations using positional encoding vectors that are added to the embedded input vectors [97]. Each embedded input element $\sigma_{i}^{(e)}$ is projected on a query vector ($q_{i}$), key vector ($k_{i}$) and a value vector ($v_{i}$) of the same dimension $d_{h}$ as the embedded input, given by:
$q_{i}=\sum_{l=1}^{d_{h}}W_{i,l}^{q}\sigma_{i,l}^{(e)}$ $k_{i}=\sum_{l=1}^{d_{h}}W_{i,l}^{k}\sigma_{i,l}^{(e)}$ $v_{i}=\sum_{l=1}^{d_{h}}W_{i,l}^{v}\sigma_{i,l}^{(e)},$ (15)
with the matrices $W^{q},W^{k}$ and $W^{v}$ to be the trainable weight matrices of dimension $d_{h}\crossproduct d_{h}$. The query, key and value matrices are then given by $Q=(q_{1},...,q_{N})$, $K=(k_{1},...,k_{N})$ and $V=(v_{1},...,v_{N})$. In multi-headed attention, each query, key and value vector is mapped to $h$ vectors with trainable weight matrices, with $h$ the number of attention heads. This is indicated in Fig. 8 (middle). A distinction has to be made between two different classes of transformers: the encoder and the decoder. Both architectures consist of the multi-head attention mechanism and a subsequent FFNN. The encoder maps an input sequence of symbolic representations to a context vector, which is used to condition the decoder. The decoder then combines this context vector with the input snapshot to generate the final output probability. Commonly, only a decoder is used for the NQS ansatz. Unlike encoders that map the input to a context vector, the decoder model is usually made autoregressive by adding a mask $M$ to the self-attention layer, which allows connections to all previous elements in the sequence but not to subsequent elements [148] and enables efficient exact sampling from the model [98, 101, 105, 103]. Then, a softmax activation function is applied to the masked dot product of the vectors $Q$ and $K$. The complete attention formalism in the decoder can be summarized by
@@
as shown in Fig. 8 (right). In Ref. [149] different variants of the attention mechanism are compared.
Transformer quantum states can learn ground state properties of various physical systems, such as the 1D transverse field Ising model and the 1D Heisenberg $J_{1}-J_{2}$ and XYZ model [98]. Comparable results to DMRG calculations have been obtained with a transformer decoder for a 1D frustrated spin model, with a relatively low number of parameters [97]. In Ref. [105], transformers have shown to be able to simulate the real-time dynamics and steady state in 1D and 2D transverse field Ising and Heisenberg models [105]. In Refs. [106, 107, 108, 109] transformers are used in the context of quantum chemistry calculations.
Small modifications of the transformer model, leading to the so-called vision transformers (ViT), inspired the use of patched transformers. This model splits the system into patches, and can be used to calculate the ground state properties of frustrated spin models [97]. A large patch size enhances the efficiency of the transformer, but on the other hand the network output dimension increases exponentially with the patch size, as it encodes the probability distribution over all possible patch states. To overcome this, large patched transformers are introduced in Ref. [101]. In this ansatz, the output of the patched transformer is passed to a patched RNN as the initial hidden state, which breaks the large inputs into smaller sub-patches, reducing the output dimension. This model has been shown to accurately capture ground state properties and phase transitions of large Rydberg systems, which can compete with quantum Monte Carlo results [101]. In Ref. [150], a combined architecture of a transformer network and a FFNN is used. Hereby, first the transformer maps the physical spins to a high-dimensional feature space. The authors argue that in this feature space the determination of the ground-state properties is simplified, requiring only a single FFNN layer with complex-valued parameters to parameterize the wave function in the second part of the network. The combined architecture is used for the ground state search of the Shastry-Sutherland Model, featuring a spin liquid phase besides phases with plaquette and antiferromagnetic order.
In Ref. [98], transformer quantum states have been used to learn ground state properties of a single system, as well as to generalize to different, unseen systems. For the latter, not only the physical degrees of freedom, but also the parameters of the Hamiltonian of the system are used as input. These parameters can be formulated as new elements that have to be passed to the embedding layer. After training the transformer quantum state for a Hamiltonian with different parameters, the transformer is able to generate the ground state for unseen Hamiltonian parameters without any additional training. Although these are with slightly larger error, more accurate results can be achieved with less training then without any a priori training. In Ref. [102] an encoder-decoder transformer is designed to learn the distribution of measurement outcomes with the Hamiltonian parameters of the system as the input. The transformer is trained on data from different interacting Rydberg arrays and is able to generalize to systems outside of the training set.
@@ 2.7.1 Fermionic Architectures
For quantum many-body systems, mostly second quantization is used, despite some exceptions e.g. for repulsively interacting, spin-polarized fermions [156], where the authors chose to model $\psi_{0}(\bm{\sigma})\mathcal{J}(\bm{\sigma})$ in Eq. (17) by a single neural network. In the works using second quantization, machine learning approaches are used to enhance the expressivity of the Slater-Jastrow ansatz. This can be done by employing NNs to parameterize the Jastrow factor $\mathcal{J}$.
One of the first examples is the RBM+PP architecture in Ref. [157], with a slightly different ansatz than Eq. (17), i.e.
@@
Currently, usually two methods are used [158]: $(i)$ The hidden fermion determinant state, where neural networks are used to replace the standard Slater determinant with a larger determinant which includes single particle orbitals from additional projected hidden fermions [159, 160]. $(ii)$ Neural backflow transformations, which add correlation by making the single particle orbitals of the Slater determinant configuration dependent, with the respective transformation learned by a neural network [161, 162, 152, 151, 139, 163]. In Ref. [158] the authors show that both $(i)$ and $(ii)$ can be written as Jastrow-like corrections to the single-particle orbitals. Furthermore, Refs. [164, 165] show that using Bloch single-particle wave functions in the Slater determinant allows to simulate large fermionic systems such as moiré materials.
@@ 2.7.2 Bosonic Architectures with Jordan Wigner Strings
The other way to simulate fermionic systems using NQS are Jordan-Wigner (JW) transformations, which are used to map the bosonic NQS to a fermionic wave function. Hence, per se bosonic architectures can be used and no special fermionic architecture is needed. The JW transformation is given by
@@
where indices refer to a one-dimensional labeling of the fermions, the $\hat{\sigma}^{\pm}$ are the spin raising and lowering operators. The resulting annihilation (creation) operators $\hat{c}^{(\dagger)}$ fulfill the fermionic commutation relations. More precisely, for two fermions at site $i$ and $j$, exchanging these fermions yields a minus sign arising from the argument of the exponential in Eq. (19). This rule does not have to be implemented in the NQS architecture itself, but only on the level of the calculation of expectation values. For an operator $\hat{O}$ with $\langle\hat{O}\rangle=\langle O^{\mathrm{loc}}_{\bm{\theta}}(\bm{\sigma})\rangle_{\bm{\sigma}}$ (see Eq. (3)) with
@@ 2.8 Other Design Choices
A second design choice is how to incorporate symmetries in the NQS training in order to restrict the optimization space to states in the target symmetry sector, hence improving the performance [62, 113, 111, 45]. Firstly, global $U(1)$ symmetries can be imposed, see e.g. Refs. [113, 111, 119], by restricting to configurations $\bm{\sigma}$ that obey the respective symmetry, e.g. magnetization or particle number conservation. For autoregressive architectures, this is done by restricting the conditional probabilities to the targeted symmetry, see e.g. Sec. 2.5.1. For non-autoregressive architectures, the Mone Carlo updates can be chosen such that all generated configurations stay in the same symmetry sector. Second, spatial symmetries can be imposed. There are different symmetrizations used in the literature [^2] We follow Ref. [45] here., which require to generate new samples $\bm{\sigma}^{S}$ that are connected by symmetry transformations $\mathcal{T}$ to the original samples $\bm{\sigma}$:
@@
    $\psi^{S}_{\bm{\theta}}(\bm{\sigma})=\sqrt{\mathrm{exp}\left(\frac{1}{|\mathcal{T}|}\sum_{\bm{\sigma}^{S}\in\mathcal{T}(\bm{\sigma})}2\mathrm{Re}\{\mathrm{log}\left[\psi_{\bm{\theta}}(\bm{\sigma}^{S})\right]\}\right)\cdot\mathrm{exp}\left(i\,\mathrm{arg}\left(\sum_{\bm{\sigma}^{S}\in\mathcal{T}(\bm{\sigma})}\mathrm{exp}\left(i\,\mathrm{Im}\{\mathrm{log}\left[\psi_{\bm{\theta}}(\bm{\sigma}^{S})\right]\}\right)\right)\right)}.$ (23)
Note that only the last keeps the autoregressive property intact since $\sum_{\bm{\sigma}^{S}\in\mathcal{T}(\bm{\sigma})}|\psi_{\bm{\theta}}^{S}(\bm{\sigma}^{S})|^{2}=\sum_{\bm{\sigma}^{S}\in\mathcal{T}(\bm{\sigma})}|\psi_{\bm{\theta}}(\bm{\sigma}^{S})|^{2}$. Furthermore, architectures that preserve certain symmetries explicitly can be used, such as group CNNs [91] and gauge equivariant neural Networks [171].
@@ 3.1.1 Variational Monte Carlo
For autoregressive architectures, $\langle\psi_{\bm{\theta}}|\psi_{\bm{\theta}}\rangle=1$ and hence the second term vanishes when calculating the derivative of Eq. (3). However, often $H_{\mathrm{loc}}(\bm{\sigma})$ is replaced by the covariate $\left(H_{\mathrm{loc}}(\bm{\sigma})-\langle H_{\mathrm{loc}}\rangle\right)$ to reduce the variance of the gradients [185, 111], leading to the same expression as Eq. (25). Another approach is to (pre-)train the NQS with experimental or numerical data, see Sec. 4.1.
@@
3.1.2 Optimization Challenges
Further improvement can be achieved using complementary optimization methods: Firstly, the NQS can be pretrained with external data, see Sec. 4. Second, the energy resulting from the VMC optimization can be improved by applying a few Lanczos steps after optimizing the network parameters using the techniques described above [204]. Applying the Hamiltonian in the Lanczos algorithm to obtain the next Krylov vector corresponds to minimizing the infidelity to the corresponding state and thus necessitates a separate training, rendering typically only very few Krylov vectors accessible. Similar in spirit to Lanczos algorithms, power methods can be used to find the ground states of (gapped) Hamiltonians [205].
@@ 3.1.2 Optimization Challenges
3.1.3 Results on the $J_{1}-J_{2}$ model
Figure 10: Comparison of energies obtained with NQS, $E_{\mathrm{NQS}}$, and matrix-product states (MPS) using density matrix renormalization group (DMRG), $E_{\mathrm{DMRG}}$ with $8192$ $SU(2)$ states [209], for ground states of the prototypical $J_{1}-J_{2}$ model at $J_{2}/J_{1}=0.5$ on $10\times 10$ lattices: We show energies obtained from FFNNs/MLPs in red (Ledinauskas et al. (2023) [206]), CNNs in blue (Liang et al. (2018) [80], Choo et al. (2019) [81], Szabo et al. (2020) [82], Liang et al. (2021) [83], Li et al. (2022) [84], Liang et al. (2022) [85], Wang et al. (2023) [86], Reh et al. (2023) [45], Chen et al. (2023) [87]), RBMs in green (Ferrari et al. (2019) [60], Nomura et al. (2021) [61], Chen et al. (2022) [204]) and transformers in yellow (Rende et al. (2023) [100]). The color coding denotes the appearance of the respective works. This plot is adapted from a table in Ref. [100].
A commonly used model for benchmarking new NQS architectures and comparing different optimization methods is the $J_{1}-J_{2}$ model,
$\mathcal{H}_{J_{1}-J_{2}}=J_{1}\sum_{\langle i,j\rangle}\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j}+J_{2}\sum_{\langle\langle i,j\rangle\rangle}\hat{\bm{S}}_{i}\cdot\hat{\bm{S}}_{j},$ (34)
with spin-$1/2$ operators $\hat{\bm{S}}_{i}$ and $\langle i,j\rangle$ ($\langle\langle i,j\rangle\rangle$) denoting nearest (next-nearest) neighbors. In this model, the nearest neighbor $J_{1}$ and the next-nearest neighbor $J_{2}$ interactions compete. For $J_{2}\to 0$ and $J_{1}>0$, the system reduces to a Heisenberg antiferromagnet, whereas for $J_{1}\to 0$ and $J_{2}>0$ the system favors antiferromagnetic stripes. In the intermediate regime around $J_{2}/J_{1}=0.5$, the nature of the ground state is not clear yet, ranging from numerical results for gapped or gapless quantum spin liquids, different types of valence bond solids or both of them [61]. With the frustration controlled by the ratio $J_{2}/J_{1}$, the $J_{1}-J_{2}$ model has become a paradigmatic model for the evaluation of the performance of NQS architectures [206, 80, 81, 82, 83, 84, 85, 86, 45, 87, 60, 61, 204, 100, 92].
The results obtained with variants of FFNNs, CNNs, RBMs and transformers for the $J_{1}-J_{2}$ model on $10\times 10$ square lattices at $J_{2}/J_{1}=0.5$ are shown in Fig. 10. It can be seen that in recent years, results from CNNs, RBMs and Transformers have become competitive with or have even outperformed DMRG results obtained for a bond dimension of $8192$ with implemented $SU(2)$ symmetry (corresponding to a bond dimension $32000$ with only $U(1)$ symmetry). In particular, recent modifications of SR as in [100, 87] have allowed to use more than $10^{5}$ parameters, systematically improving results obtained with smaller NQS with around $10^{3}-10^{4}$ parameters. However, also for $10^{3}-10^{4}$ parameters some works have obtained energy errors that are competitive with the NQS architectures using more than one order of magnitude more parameters [61, 86, 45]. It hence becomes evident that in principle many NQS architectures can achieve competitive results to conventional methods, but details on the implementation and the optimization procedure can have a significant impact on the performance. Consequently, architecture, its hyperparameters and optimization parameters have to be carefully chosen, but at the same time they can mostly only be determined by try and error.
@@ 3.2 Excited States
The former usually rely on the same VMC scheme, but with a restriction to the targeted symmetry sector implemented in the wave function or the training loss [53, 35, 132, 62, 170, 210]. This enables e.g. the calculation of quasiparticle dispersions from NQS [53, 35, 132, 62, 69]:
In Ref. [62], excited states are targeted using quantum-number projections [211], i.e. for a NQS parameterization $\psi_{\bm{\theta}}(\bm{\sigma})$ one defines the wave function using the total momentum $\bm{k}$ projection
$\psi_{\bm{k}}(\bm{\sigma})=\sum_{\bm{R}}e^{-i\bm{k}\cdot\bm{R}}\psi_{\bm{\theta}}(\hat{T}_{\mathrm{R}}\bm{\sigma})$ (35)
with the translation operator $\hat{T}_{\bm{R}}$, shifting the system by the vector $\bm{R}$. This can be combined with other quantum-number projections, including e.g. spin parity and spatial symmetries, to improve the accuracy. In another approach [35], specific momentum $\bm{k}_{\mathrm{target}}$ sectors are targeted by adding a mean square error between $\bm{k}_{\mathrm{target}}$ and $\bm{k}_{\mathrm{NQS}}$, see Fig. 11a. Hereby, the momentum of the NQS is given by